You are seeking $1.5 million from a venture capitalist to finance the launch of your online financial search engine. You and the VC agree that your venture is currently worth$3 million, and when the company goes public in an IPO in five years, it is expected to have a market capitalization of $20 million. Given the company’s stage of development, the VC requires a 50 percent return on investment. 

What fraction of the firm will the VC receive in exchange for its $1.5 million investment in your company?


1
Expert's answer
2021-10-28T16:23:42-0400

Current VC investment =$1.5 million

Consent valuation = $3 million

Expected valuation after 5 years = $20 million

After 5 years, VC investment value will be:

Let the value of VC be "x"


"x =1.5(1+50\%)^5 =\$11.3906"

Fraction of the firm the VC will receive is


"\frac{x}{20 million}=\frac{11.3906}{20}=56.953\%"